# Adaptive Incentive Design with Learning Agents

### v2

@@ Abstract
How can the system operator learn an incentive mechanism that achieves social optimality based on limited information about the agents’ behavior, who are dynamically updating their strategies? To answer this question, we propose an adaptive incentive mechanism. This mechanism updates the incentives of agents based on the feedback of each agent’s externality, evaluated as the difference between the player’s marginal cost and society’s marginal cost at each time step. The proposed mechanism updates the incentives on a slower timescale compared to the agents’ learning dynamics, resulting in a two-timescale coupled dynamical system. Notably, this mechanism is agnostic to the specific learning dynamics used by agents to update their strategies. We show that any fixed point of this adaptive incentive mechanism corresponds to the optimal incentive mechanism, ensuring that the Nash equilibrium coincides with the socially optimal strategy. Additionally, we provide sufficient conditions that guarantee the convergence of the adaptive incentive mechanism to a fixed point. Our results apply to both atomic and non-atomic games. To demonstrate the effectiveness of our proposed mechanism, we verify the convergence conditions in two practically relevant games: atomic networked quadratic aggregative games and non-atomic network routing games.
@@ 1 Introduction
Incentive mechanisms play a crucial role in many societal systems, where the outcomes are governed by the selfish interactions of a large number users (or algorithms on their behalf) and the system operator. The outcome of such strategic interactions, characterized by the Nash equilibrium, is often sub-optimal because of the fact that individual agents often do not account for the externality of their actions (i.e., how their actions affect the costs of others), when minimizing their own costs. An important way to address the sub-optimality is to provide agents with incentives that align their individual goal of cost minimization with the goal of minimizing the total cost of the societal system[Levin, 1985, Ho et al., 1982, Başar, 1984].
The system operator faces two key challenges when designing such incentive mechanisms. First, due to the system’s scale and privacy concerns, the designer often has limited information about individual agents’ costs and the cost function of the entire system. Second, the system operator often needs to consider the learning behavior of agents, who repeatedly update their strategies in response to the incentive mechanism, especially when the physical system experiences a random shock and players are trying to reach a new equilibrium [Barrera and Garcia, 2014, Como and Maggistro, 2021, Maheshwari et al., 2021]. As a result, the designer cannot infer the equilibrium to predict agents’ behavior, making incentive design a challenging task. This leads to the following question:
How can the system operator learn an incentive mechanism that achieves social optimality based on limited information about the behavior of agents who are adaptively updating their strategies?
@@
To answer this question, we propose an adaptive incentive mechanism that updates the incentives based on the behavior of agents, which is repeatedly updated as part of some learning process, and the gradient information of the cost functions. This results in a coupled dynamical system comprising incentive and strategy updates. Our proposed framework can be applied to both atomic and non-atomic games.
@@
Our proposed incentive mechanism has three key features. First, the incentive update incorporates the externality incurred by the agents’ current strategy, which is evaluated as the difference between their own marginal cost and the marginal cost for the entire system. This feature guarantees that any fixed point of the coupled learning dynamics induces a socially optimal outcome: the incentive provided to each agent equals their externality given their equilibrium strategy, and the associated Nash equilibrium is a socially optimal strategy (Proposition 3.1). Second, the incentive mechanism only requires limited information about the behavior of agents. Specifically, it is agnostic to the strategy update dynamics used by agents and only requires information about the gradient of the cost function of agents at the current strategy to evaluate the externality (Remark 2.2). Third, the incentive update occurs at a slower timescale compared to the agents’ strategy updates. This slower evolution of incentives is a desirable characteristic because frequent incentive updates may hamper agents’ participation. It allows agents to consider the incentives as static while updating their strategies.
@@
We provide sufficient conditions on the underlying game such that the set of fixed points is a singleton, making the socially optimal incentive mechanism unique (Proposition 3.1). We also characterize sufficient conditions that guarantee the convergence of the coupled dynamical system (Proposition 3.3). Since the convergent strategy profile and incentive mechanism correspond to a fixed point that is also socially optimal, these sufficient conditions ensure that the dynamic incentive mechanism eventually induces a socially optimal outcome in the long run. To prove this result, we leverage the timescale separation between the strategy update and the incentive update, using results from the theory of two-timescale dynamical systems [Borkar, 1997]. This theory allows us to analyze the convergence of the coupled dynamical system in two steps by decoupling the convergence of the strategy update and the incentive mechanism update. First, the convergence of the strategy update (on the faster timescale) is analyzed by viewing the incentive mechanism (on the slower timescale) as static. In our work, we off-load this convergence to the rich literature on learning in games (e.g., [Fudenberg and Levine, 1998, Sandholm, 2010, Monderer and Shapley, 1996]). Second, the convergence of the incentive mechanism update is analyzed via the associated continuous-time dynamical system, evaluated at the converged value of the strategy update (i.e., Nash equilibrium). The sufficient conditions developed in Proposition 3.3 are built on results on the convergence of non-linear dynamical systems from cooperative dynamical systems theory [Hirsch, 1985] and Lyapunov-based methods [Sastry, 2013].
@@
To demonstrate the usefulness of the adaptive incentive mechanism, we apply it to two classes of practically relevant games: (i) atomic networked quadratic games and (ii) non-atomic routing games. In atomic networked quadratic games, each agent’s cost function depends on their own strategy as well as the aggregate strategies of their opponents. This aggregation is done through a linear combination of the neighboring agents’ strategies with weights characterized by a network matrix. Our proposed incentive mechanism enables the system operator to adaptively update incentives, based on the externality of each agent’s strategy on their neighbors, while the agents are learning the equilibrium strategies. Our results when specialized to this setup provide sufficient conditions on the network matrix to ensure that the system converges to a socially optimal outcome.
@@
Furthermore, in the non-atomic routing games, agents (travelers) make routing decisions in a congested network with multiple origin-destination pairs. The system operator imposes incentives in the form of toll prices on each network edge. Our proposed incentive mechanism adaptively gets updated using only the observed flow on edges and the gradient of the edge latency function. Agents can adopt a variety of strategy updates that lead to the equilibrium of the routing game. We show that the adaptive incentive mechanism converges to the toll price that minimizes total congestion.
The article is organized as follows: in Sec. 2, we describe the setup of both atomic and non-atomic game considered here. In addition, we also provide the joint strategy and incentive update considered in this paper. We present the main results on the fixed point and convergence of adaptive incentive mechanism in Sec. 3. In Sec. 4, we study the application of our proposed adaptive incentive mechanism on the atomic networked aggregative games (Sec. 4.1) and non-atomic routing game (Sec. 4.2). We conclude our work in Sec. 5.
@@ 1.1 Related Works
Two-timescale Learning Dynamics: Learning dynamics where incentives are updated at a slower timescale than the strategy of agents have been studied in [Mojica-Nava et al., 2022, Chen et al., 2023, Poveda et al., 2017, Ochoa and Poveda, 2022, Ratliff and Fiez, 2020, Liu et al., 2021, Li et al., 2023b]. Specifically, [Mojica-Nava et al., 2022] studies Stackelberg games with a single leader and a population of followers. In their approach the leader deploys gradient-based update and followers update their strategies using replicator dynamics. [Chen et al., 2023] and [Poveda et al., 2017] consider incentive design in affine congestion games, where the incentives are updated using a distributed version of gradient descent. [Ochoa and Poveda, 2022] studies incentive design to control traffic congestion on a single highway using gradient based incentive updates. [Ratliff and Fiez, 2020] studies the problem of incentive design while learning the cost functions of agents. The authors assume that both the cost functions and incentive policies are linearly parameterized, and the incentive updates rely on the knowledge of agents’ strategy update rules instead of just the current strategy as in our setting. Additionally, [Liu et al., 2021, Li et al., 2023b] propose a two-timescale discrete-time learning dynamics, where agents adopt a mirror descent-based strategy update, and the social planner updates the incentive parameter according to a gradient based method. The convergence of these gradient-based learning dynamics relies on the assumption that the social cost given agents’ equilibrium strategy is convex in the incentive parameter. However, the convexity assumption is restrictive as highlighted by an example in Appendix A.
@@
Same-timescale Learning Dynamics: The problem of steering non-cooperative agents to a desired Nash equilibrium by considering an incentive update which runs on the same timescale as strategy updates is studied in [Shakarami et al., 2022a, Shakarami et al., 2022b, Zhang et al., 2023]. Particularly, in [Shakarami et al., 2022a] the authors study such updates in a specific setting of quadratic aggregative games. In [Shakarami et al., 2022b], the authors consider a setup where the agents’ cost is only dependent on their own action and the price signal which is provided by operator which is used by agents based on their individual reliability over the operator. The aim is to design this price signal so that the reliability on the price signal is retained while steering the equilibrium to the desired point. In [Zhang et al., 2023], the authors consider the problem of steering no-regret learning agents to optimal equilibrium but the computation of incentive mechanism requires solving an optimization problem at each time step.
@@
Learning in Stackelberg Games: Our work is also related to the literature on learning in Stackelberg games, where the planner often has limited information about the game between agents and need to design optimal mechanism by dynamically receiving feedback from agents’ responses (for instance, refer to [Blum et al., 2014, Letchford et al., 2009, Peng et al., 2019, Bai et al., 2021, Cui et al., 2024, Jalota et al., 2022, Maheshwari et al., 2023]). Typically, this line of research posit structural assumptions on the game between followers (e.g., finite action space of agents, or linearly parametrized utility function of agents)[Blum et al., 2014, Letchford et al., 2009, Peng et al., 2019, Bai et al., 2021, Cui et al., 2024, Jalota et al., 2022], or only ensures convergence to a locally optimal solution [Maheshwari et al., 2023].
In comparison to the preceding three lines of research, our approach of externality-based incentive design is applicable for a broad class of games (with continuous action spaces and potentially non-linear utility functions), while providing sufficient conditions to ensure asymptotic convergence to the optimal solution.
@@
Given a function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$, we use $\nabla_{x_{i}}f(x)$ to denote $\frac{\partial f}{\partial x_{i}}(x)$, the derivative of $f$ with respect to $x_{i}$ for any $i\in\{1,2,...,n\}$ and $\nabla f$ to be the Jacobian of the function. For any set $A$ we use $\textsf{conv}(A)$ to denote the convex hull of the set. For any set $X\subseteq\mathbb{R}^{n}$, we call a function $f:X\rightarrow\mathbb{R}$ to be Lipschitz if there exists a positive scalar $L$ such that $\|f(x)-f(x^{\prime})\|\leqslant L\|x-x^{\prime}\|$ for every $x,x^{\prime}\in X$. For any vector $x\in X$ and any positive scalar $r>0$, the set $\mathcal{B}_{r}(x)=\{x^{\prime}\in X|\|x^{\prime}-x\|<r\}$ denotes the $r$-radius neighborhood of the vector $x$. For any set $X$, we define $\textsf{boundary}(X)$ and $\textsf{int}(X)$ to be the boundary and interior of set $X$, respectively. Finally, for any function $f(\cdot)$, we denote the domain of the function by $\textsf{dom}(f)$. For any vector $x\in\mathbb{R}^{n}$, we define $\textsf{diag}(x)\in\mathbb{R}^{n\times n}$ to be a diagonal matrix with diagonal entries corresponding to $x$.
@@ 2 Model
We introduce both atomic and non-atomic static games in Sec. 2.1. In Sec. 2.2, we present the adaptive incentive design approach proposed in this work.
@@ 2.1.1 Atomic Games
Consider a game $G$ with a finite set of players $\mathcal{I}$. The strategy of each player $i\in\mathcal{I}$ is ${x}_{i}\in X_{i}$, where $X_{i}$ is a non-empty and closed interval in $\mathbb{R}$. The joint strategy profile of all players is ${x}=({x}_{i})_{i\in\mathcal{I}}$, and the set of all joint strategy profiles is $X\coloneqq\prod_{i\in\mathcal{I}}X_{i}$. The cost function of each player $i\in\mathcal{I}$ is $\ell_{i}:\mathbb{R}^{|\mathcal{I}|}\to\mathbb{R}$ [^1] We measure the outcome of our games by costs instead of utilities. Equivalently, the utility of each player is the negative value of the cost.
A social planner designs incentives by setting a payment $p_{i}x_{i}\in\mathbb{R}$ for each player $i$ that is linear in their strategy $x_{i}$. Here, $p_{i}\in\mathbb{R}$ represents the marginal payment for every unit increase in strategy of player $i$. The value of $p_{i}x_{i}$ can either be negative or positive, which represents a marginal subsidy or a marginal tax, respectively. Given the incentive mechanism $p=(p_{i})_{i\in I}$, the total cost of each player $i\in\mathcal{I}$ is:
@@
The cost function $\ell_{i}({x}_{i},{x}_{-i})$ is strongly convex in ${x}_{i}$ for all $i\in\mathcal{I}$ and all ${x}_{-i}=({x}_{j})_{j\in\mathcal{I}\setminus\{i\}}$. Moreover, $\ell_{i}(\cdot)$ is continuously differentiable and its gradient is Lipschitz. Additionally, the social cost function $\Phi(x)$ is twice continuously differentiable, has Lipschitz gradient, and is strictly convex.
Under Assumption 2.1, it is known that, for any $p\in\mathbb{R}^{|\mathcal{I}|}$, the Nash equilibrium $x^{\ast}(p)$ exists, is unique and is Lipscthiz continuous in $p$ [Dafermos, 1988]. Moreover, the socially optimal strategy $x^{\dagger}$ is also unique.

### v3

@@ Abstract
We propose an adaptive incentive mechanism that learns the optimal incentives in environments where players continuously update their strategies. Our mechanism updates incentives based on each player’s externality, defined as the difference between the player’s marginal cost and the operator’s marginal cost at each time step. The proposed mechanism updates the incentives on a slower timescale compared to the players’ learning dynamics, resulting in a two-timescale coupled dynamical system. Notably, this mechanism is agnostic to the specific learning dynamics used by players to update their strategies. We show that any fixed point of this adaptive incentive mechanism corresponds to the optimal incentive mechanism, ensuring that the Nash equilibrium coincides with the socially optimal strategy. Additionally, we provide sufficient conditions under which the adaptive mechanism converges to a fixed point. Our results apply to both atomic and non-atomic games. To demonstrate the effectiveness of our proposed mechanism, we verify the convergence conditions in two practically relevant classes of games: atomic aggregative games and non-atomic routing games.
@@ 1 Introduction
Incentive mechanisms play a crucial role in many societal systems, where outcomes are governed by the interactions of a large number of self-interested users (or algorithms acting on their behalf). The outcome of such strategic interactions, characterized by the Nash equilibrium, is often suboptimal because individual players typically do not account for the externality of their actions (i.e., how their actions affect the costs of others) when minimizing their own costs. An important way to address this suboptimality is to provide players with incentives that align their individual goal of cost minimization with the goal of minimizing the total cost of the societal system [Levin, 1985, Başar, 1984]. However, this problem becomes more challenging as the system operator often needs to account for the learning behavior of players, who repeatedly update their strategies in response to the incentive mechanism, especially when the physical system experiences a random shock and players are learning to reach a new equilibrium [Barrera and Garcia, 2014, Como and Maggistro, 2021, Maheshwari et al., 2021].
@@
To address this challenge, we propose an adaptive incentive mechanism that adjusts incentives based on the strategies of players, who repeatedly update their strategies as part of a learning process. This results in a coupled dynamical system that comprises both incentive and strategy updates.
@@
Our proposed incentive mechanism has four key features. Firstly, our framework applies to both atomic and non-atomic games. Secondly, the incentive update incorporates the externality generated by the players’ current strategies, quantified as the difference between their own marginal cost and the marginal cost for the entire system. Thirdly, the incentive mechanism is agnostic to the strategy update dynamics used by players and requires only oracle access to either the gradient (in atomic games) or the value (in non-atomic games) of the cost function, given the current strategy, to evaluate the externality. Finally, the incentive update occurs on a slower timescale compared to the players’ strategy updates. This slower evolution of incentives is a desirable characteristic because frequent incentive updates often hinder players’ participation.
We prove that any fixed point of the coupled incentive and strategy updates leads to a socially optimal outcome. Specifically, at any fixed point, the incentive provided to each player equals the externality of the equilibrium strategy, ensuring that the resulting Nash equilibrium is socially optimal (Proposition 3.1). Additionally, we establish sufficient conditions on the underlying game that ensure the fixed point – coinciding with the socially optimal incentive mechanism – is unique (Proposition 3.1).
@@
We characterize sufficient conditions for (both local and global) convergence of the coupled dynamical system to the fixed points (Proposition 3.3). Since the convergent strategy profile and incentive mechanism correspond to a socially optimal outcome, these sufficient conditions ensure that the coupled dynamical system induce a socially optimal outcome in the long run. Our analysis builds on the theory of two-timescale dynamical systems [Borkar, 1997]. Due to the timescale separation between the strategy and incentive updates, we can decouple the convergence of the strategy update from that of the incentive update. First, the convergence of the strategy update, which evolves on the faster timescale, is analyzed by treating the incentive mechanism, which evolves on the slower timescale, as static. In our work, we offload this analysis to the extensive literature on learning in games (e.g., [Mertikopoulos et al., 2024, Leslie and Collins, 2006, Benaïm et al., 2005, Sandholm, 2010, Swenson et al., 2018]). Second, the convergence of the incentive mechanism update is examined through the corresponding continuous-time dynamical system, evaluated at the fixed point of the strategy update (i.e., the Nash equilibrium).
@@
To demonstrate the usefulness of the adaptive incentive mechanism, we apply it to two practically relevant classes of games: (i) atomic aggregative games and (ii) non-atomic routing games. In atomic aggregative games, each player’s cost function depends on their own strategy as well as the aggregate strategies of their opponents. This aggregation is performed through a linear combination of neighboring players’ strategies, with weights characterized by a network matrix. Our proposed incentive mechanism enables the system operator to adaptively adjust incentives based on each player’s externality on their neighbors while players learn their equilibrium strategies. When applied to this setting, our results provide sufficient conditions on the network matrix to ensure global convergence to a socially optimal outcome.
@@
Furthermore, in non-atomic routing games, players (travelers) make routing decisions in a congested network with multiple origin-destination pairs. The system operator imposes incentives in the form of toll prices on network edges. Our proposed incentive mechanism is adaptively updated based solely on the observed edge flows and the gradient of the edge latency functions. Players can follow various strategy update rules that lead to the equilibrium of the routing game. We show that the adaptive incentive mechanism locally converges to the toll prices that minimize total congestion.
The article is organized as follows: In Sec. 2, we describe the setup for both atomic and non-atomic games and introduce the joint strategy and incentive update framework. Sec. 3 presents our results on the fixed points being socially optimal, and sufficient conditions for local and global convergence in general games. In Sec. 4, we apply these convergence results to atomic aggregative games (Sec. 4.1) and non-atomic routing games (Sec. 4.2). Finally, we conclude in Sec. 5.
@@ 1.1 Related Works
Two-timescale Learning Dynamics: Learning dynamics in which incentives are updated on a slower timescale than players’ strategies have been studied in [Mojica-Nava et al., 2022, Chen et al., 2023, Poveda et al., 2017, Ochoa and Poveda, 2022, Liu et al., 2022, Li et al., 2023, Alpcan and Pavel, 2009, Alpcan et al., 2009]. Specifically, [Mojica-Nava et al., 2022] examines Stackelberg games with a single leader and a population of followers, where the leader employs gradient-based updates while the followers adjust their strategies using replicator dynamics. Moreover, [Chen et al., 2023, Poveda et al., 2017] focus on incentive design in affine congestion games, where incentives are updated using a distributed version of gradient descent. Similarly, [Ochoa and Poveda, 2022] studies incentive design for traffic control on a single highway through gradient-based incentive updates. Additionally, [Liu et al., 2022, Li et al., 2023] propose a two-timescale discrete-time learning dynamic in which players update their strategies using mirror descent, while the system operator adjusts the incentive parameter via a gradient-based method. Furthermore, [Alpcan and Pavel, 2009, Alpcan et al., 2009] study the convergence of gradient-based incentive updates when the system operator has access to the gradient of the equilibrium strategy with respect to the incentive.
All of these works adopt gradient-based incentive updates. In such approaches, ensuring that the fixed point is socially optimal relies on the assumption that the equilibrium social cost is a convex function of the incentive parameter [Mojica-Nava et al., 2022, Chen et al., 2023, Poveda et al., 2017, Ochoa and Poveda, 2022, Liu et al., 2022, Li et al., 2023] or that the gradient of the equilibrium strategy with respect to the incentive is non-singular [Alpcan et al., 2009, Alpcan and Pavel, 2009]. However, these assumptions are restrictive and often do not hold, even in simple games. In Appendix A, we provide a counterexample—a two-link routing game—in which both the convexity and non-singular gradient assumptions fail to hold.
@@
Single-timescale Learning Dynamics: The problem of steering non-cooperative players toward a desired Nash equilibrium using an incentive update that operates on the same timescale as strategy updates has been studied in [Shakarami et al., 2023, Shakarami et al., 2022, Zhang et al., 2023, Ratliff and Fiez, 2020]. Specifically, [Shakarami et al., 2023] examines such updates in the setting of quadratic aggregative games. In [Shakarami et al., 2022], the authors consider a scenario where players’ costs depend only on their own actions and a price signal provided by an operator. In [Zhang et al., 2023], the authors address the problem of guiding no-regret learning players toward an optimal equilibrium; however, their approach requires solving an optimization problem at each time step to compute the incentive mechanism. The work in [Ratliff and Fiez, 2020] explores incentive design while simultaneously learning players’ cost functions. The authors assume that both cost functions and incentive policies are linearly parameterized, with incentive updates relying on knowledge of the players’ strategy update rules rather than solely on their current strategies, as in our setting.
@@
Learning in Stackelberg Games: Our work is also related to the literature on learning in Stackelberg games, where the planner often has limited information about the interactions between players and must design an optimal mechanism by dynamically incorporating feedback from players’ responses (see, e.g., [Blum et al., 2014, Letchford et al., 2009, Peng et al., 2019, Bai et al., 2021, Cui et al., 2024, Jalota et al., 2022, Maheshwari et al., 2023]). This line of research typically imposes structural assumptions on the game among followers, such as a finite action space or linearly parameterized utility functions [Blum et al., 2014, Letchford et al., 2009, Peng et al., 2019, Bai et al., 2021, Cui et al., 2024, Jalota et al., 2022]. Alternatively, some works, such as [Maheshwari et al., 2023], focus on ensuring convergence only to a locally optimal solution.
Compared to the preceding three lines of research, we introduce a novel externality-based adaptive incentive design that applies to both atomic and nonatomic games, accommodates continuous action spaces, and allows for nonlinear utility functions. Unlike gradient-based incentive updates, externality-based updates ensure that any fixed point of the dynamics is socially optimal without requiring the equilibrium social cost function to be convex in the incentive vector or the gradient of equilibrium strategy with respect to the incentive to be non-singular. Furthermore, our incentive update is agnostic to the players’ learning dynamics and relies only on oracle access to zeroth-order or first-order information about players’ costs given their current strategies.
@@
Given a function $f:\mathbb{R}^{n}\rightarrow\mathbb{R}$, we use $\nabla_{x_{i}}f(x)$ to denote the partial derivative of $f$ with respect to $x_{i}$ for any $i\in\{1,2,...,n\},$ and $\nabla f(x)$ to denote the gradient of the function. For any set $A$, we use $\textsf{conv}(A)$ to denote its convex hull. For any set $X\subseteq\mathbb{R}^{n}$, a function $f:X\rightarrow\mathbb{R}$ is Lipschitz if there exists a positive scalar $L$ such that $\|f(x)-f(x^{\prime})\|\leqslant L\|x-x^{\prime}\|,$ for every $x,x^{\prime}\in X$. For any vector $x\in X$ and any positive scalar $r>0$, the set $\mathcal{B}_{r}(x)=\{x^{\prime}\in X|\|x^{\prime}-x\|<r\}$ denotes the $r$-radius neighborhood of the vector $x$. For any set $X$, we define $\textsf{boundary}(X)$ and $\textsf{int}(X)$ to be the boundary and interior of set $X$, respectively. Finally, for any function $f(\cdot)$, we denote the domain of the function by $\textsf{dom}(f)$. For any vector $x\in\mathbb{R}^{n}$, we define $\textsf{diag}(x)\in\mathbb{R}^{n\times n}$ to be a diagonal matrix with diagonal entries corresponding to $x$.
@@ 2 Model
We introduce both atomic and non-atomic static games in Sec. 2.1. In Sec. 2.2, we present our proposed adaptive incentive design approach.
@@ 2.1.1 Atomic Games
Consider a game $G$ with a finite set of players $\mathcal{I}$. The strategy of each player $i\in\mathcal{I}$ is denoted by ${x}_{i}\in X_{i}$, where $X_{i}$ is a non-empty, closed interval in $\mathbb{R}$. The joint strategy profile of all players is given by ${x}=({x}_{i})_{i\in\mathcal{I}}$, and the set of all joint strategy profiles is $X\coloneqq\prod_{i\in\mathcal{I}}X_{i}$. The cost function of each player $i\in\mathcal{I}$ is represented as $\ell_{i}:\mathbb{R}^{|\mathcal{I}|}\to\mathbb{R}$.
A system operator designs incentives by setting a payment $p_{i}x_{i}\in\mathbb{R}$ for each player $i$, which is linear in their strategy $x_{i}$. Here, $p_{i}\in\mathbb{R}$ represents the marginal payment for every unit increase in the strategy of player $i$. The value of $p_{i}x_{i}$ can be either negative or positive, representing a marginal subsidy or a marginal tax, respectively. Given the incentive mechanism $p=(p_{i})_{i\in I}$, the total cost for player $i\in\mathcal{I}$ is:
@@
For any $p\in\mathbb{R}^{|\mathcal{I}|}$, the Nash equilibrium $x^{\ast}(p)$ is unique and Lipschitz continuous in $p$. Moreover, the social cost function $\Phi(x)$ is continuously differentiable, has a Lipschitz gradient, and is strictly convex in $x$.
Assumption 2.1 is widely adopted in the literature to study incentive design in atomic games (e.g., [Li et al., 2024, Liu et al., 2022, Li et al., 2023, Shakarami et al., 2023]), either directly or through other conditions that guarantee this [^1] Uniqueness and Lipschitz continuity of $x^{\ast}(p)$ hold if, for every $i\in\mathcal{I}$ and ${x}_{-i}=({x}_{j})_{j\in\mathcal{I}\setminus\{i\}}$, the cost function $\ell_{i}({x}_{i},{x}_{-i})$ is strongly convex in ${x}_{i}$ and $\ell_{i}(\cdot)$ is continuously differentiable with a Lipschitz gradient [Dafermos, 1988]..